OpenStudy (anonymous):

4. Which of these groups of side lengths fit the Pythagorean Theorem? 8, 6, 10 6, 8, 12 5 ,4, 3 7, 8, 9

OpenStudy (anonymous):

The first one

OpenStudy (anonymous):

How come it is the first?

OpenStudy (anonymous):

well you would need to go through each problem and find if it makes the following equation true a^2+b^2=c^2

OpenStudy (anonymous):

a^2 + b^2 = c^2 If you replace the variables with those numbers you will see :)

OpenStudy (anonymous):

the first one is correct because 8^2+6^2 is 100 10^2 =100 so a^2+b^2=c^2 works

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

you are welcome